# A Quadratic Programming Approach to Flight Envelope Protection Using Control Barrier Functions

### v1

@@ 1 Introduction
Recent advancements in control theory offer an alternative perspective on designing FEP systems, framing the problem as a closed-loop state-constrained control problem. In such a framework, the controllers not only track desired set points but also dynamically enforce safety constraints throughout the system’s operation, both during steady-state and transient phases. Control Barrier Functions (CBFs) have emerged as a powerful tool within this class of safety-critical control methods. CBFs guarantee the forward invariance of a defined safe set, providing formal assurances that the system remains within prescribed limits under all operational conditions [11]. The optimization-based nature of CBFs allows for seamless integration of a wide range of constraints, including physical limitations such as load factor and stall prevention, as well as operational constraints like altitude and obstacle avoidance [12, 13]. A significant advantage of CBF-based methods is their ability to augment existing control architectures. Rather than requiring a complete redesign, CBFs can function as an additional safety layer, leveraging the existing closed-loop dynamics as part of the optimization problem [14, 15]. This compatibility facilitates their implementation in modern aerospace systems, while the real-time adaptability of CBFs enables optimal performance under varying conditions, disturbances, and uncertainties [16, 11, 17]. Despite their potential, the application of CBFs to FEP remains relatively underexplored in the flight control community.
@@ 2.1 The Mathematical Concept of Control Barrier Functions
The above definitions lead to a less restrictive version of Eq. (7) as it weakens the requirement to that in Eq. (8). The derived concept is illustrated in Fig. 1. By definition, a valid barrier function $h(x(t))$ is defined as positive within the safe set $S$. To provide safety guarantees and forward invariance within $S$, the condition on the boundary of the safety set needs to be $\dot{h}\geq 0$. When this condition is satisfied, it can be proven that the system remains safe.
@@
Since the relevant nonlinear dynamics of the flight systems considered in this work are of a relative degree greater than one, we extend the discussion of CBFs with the notion of High Order Barrier Functions (HOBFs) and High Order Control Barrier Functions (HOCBFs), as proposed in [22]. The control-theoretical construct of CBFs, introduced here, along with its extension to HOCBFs, provides a unified methodology for enforcing state constraints derived from operational limits, such as aerodynamic, structural, and performance constraints, which are critical for ensuring safe flight in aerospace systems with complex and nonlinear behaviors. By establishing these theoretical foundations, the framework sets the stage for its application to practical systems, as demonstrated in this work through its integration into an FEP scheme. we consider a time-varying function to define an invariant set for system (1). For a $d$-th order differentiable function $h:\mathbb{R}^{n}\times[t_{0},\infty)\to\mathbb{R}$ (where $t_{0}$ denotes the initial time), we define a series of functions $\psi_{0},\psi_{1},\ldots,\psi_{d}$ recursively as:
@@
When a control-affine of the form (9) is considered, HOBFs can be extended to High Order Control Barrier Functions (HOCBFs) by incorporating control inputs.
@@
This framework introduces a systematic approach to ensure system safety by directly incorporating control inputs into the analysis. Unlike conventional reference filtering methods, HOCBFs enforce safety constraints as an inherent property of the system dynamics, ensuring that operational limits—such as those on angle of attack, load factor, or aerodynamic stability—are satisfied at all times. This makes HOCBFs particularly well-suited for applications in aerospace systems, where higher-order dynamics, such as second-order attitude constraints, must be carefully managed to guarantee safe and efficient flight operations.
@@ 2.2 Considered Nonlinear Missile Dynamics
Under the assumption that the roll motion is rapidly stabilized, the missile motions can be decoupled into two perpendicular and identical channels (i.e., pitch and yaw). Hence, the pitch motion is only considered as follows:
@@
$\dot{q}$ $=\frac{QSd}{I_{yy}}\left(C_{M,\alpha}(M)\alpha(t)+C_{M,q}(M)\frac{d}{2V}q(t)+C_{M,\delta}(M)\delta(t)\right),$ (22)
$n_{z}$ $=\frac{QS}{mg}\left(C_{Z,\alpha}(M)\alpha(t)+C_{Z,\delta}(M)\delta(t)\right),$ (23)
@@
3 The Flight Envelope Protection Problem
For flight systems, aerodynamics are the primary driver of the vehicle’s dynamics, governing its performance and stability during flight. Aerodynamic forces and moments arise from the interaction between the vehicle and the surrounding air, affected by control surfaces and incidence angles such as the angle of attack $\alpha(t)$ and sideslip angle $\beta(t)$. A flight envelope, for any flight vehicle, can be effectively defined and graphically illustrated using a V-n diagram, as shown in Fig. 3. This diagram illustrates the upper and lower load limits corresponding to different airspeeds, $V$. These limits provide critical insights into the vehicle’s operational constraints and form the basis for understanding load restrictions.
$x(t_{0})\in S(t)\implies x(t)\in S(t),\quad\forall t\geq t_{0},$ (26)
4 Nonlinear Flight Control System Design with Guaranteed FEP
This section presents the design of a nonlinear flight control system integrated with a FEP safety filter. First, a cascaded nonlinear attitude flight controller is designed to regulate the nonlinear longitudinal missile dynamics. Following this, a CBF-based safety filter is proposed to ensure adherence to predefined safety constraints, effectively maintaining the system’s operation within the defined flight envelope while minimally influencing control performance when the dynamics are far from the envelope boundaries.
4.1 Nonlinear Attitude Controller Design
$\dot{\alpha}_{c}(t)=K_{\alpha}(\alpha_{r}(t)-\alpha(t)),$ (27)
$\dot{q}_{c}(t)=K_{q}(q_{r}(t)-q(t)),$ (29)
Controller Gain Conservative Aggressive
$K_{\alpha}$ 4.00 36.53
$K_{q}$ 1.2 12.22
Remark 2
It should be noted that, as will be shown in the following section, the potentially unsafe control surface deflection $\delta^{*}(t)$ will be adjusted to ensure that the flight vehicle remains within its designated flight envelope. However, this adjustment comes at the cost of deviating from the reference command $\alpha_{r}(t)$. This deviation can introduce challenges, particularly when integral controllers are employed within the integrated flight control system to ensure adequate tracking performance. In such cases, integrator windup may occur, leading to undesirable control behavior and degraded system performance. To mitigate these effects, an integrator reset logic must be implemented to prevent excessive error accumulation and maintain stable control performance.
4.2 CBF-Based Nonlinear Safety Filter Design
Fig. 6 illustrates the qualitative relationship between airspeed $V$ and the permissible limits on the angle of attack $\alpha(t)$, which must not be violated to ensure safe operation. The dashed-dotted black lines represent structural constraints, the dotted black lines indicate stall boundaries, and the solid red lines delineate the combined overall operational limits of $\alpha(t)$ as a function of airspeed. The operational limits on $\alpha(t)$ imposed by structural load constraints can be derived from Eq. (23), considering the strictly defined structural load factor limits $n_{z,\text{max,structural}}$ and $n_{z,\text{min,structural}}$ the equation can be inverted to obtain the corresponding maximum and minimum permissible angles of attack $\alpha_{\text{max,structural}}$ and $\alpha_{\text{min,structural}}$. The structural constraints thereby impose an inverse-square dependency of $\alpha(t)$ on $V$, owing to the quadratic dependence of $Q$ on $V$. In contrast, stall limits are typically assumed to be independent of airspeed, meaning that the maximum and minimum allowable angles of attack due to stall, $\alpha_{\text{max,stall}}$ and $\alpha_{\text{min,stall}}$, remain constant across the flight envelope. As a result, the operational bounds on $\alpha(t)$ are shaped by the interplay between structural and stall constraints. At lower airspeeds, stall limits dominate the constraints on $\alpha(t)$. In contrast, at higher airspeeds, the range of feasible $\alpha(t)$ values is increasingly restricted by structural considerations, as illustrated in Fig. 6 with solid red lines. By considering the described two origins of constraints on $\alpha$, the overall operational limits can be formally expressed as:
where:
$\alpha_{\text{crit,min}}$ $=\max(\alpha_{\text{min,stall}}(t),\alpha_{\text{min,structural}}(t)).$
Figure 7: Block diagram of control architecture with the integrated QP-based safety filter for FEP.
To dynamically enforce the derived angle-of-attack constraints, a HOCBF-based safety filter is implemented. The realized safety filter integration with the previously established flight control system is illustrated in a block diagram in Fig. 7. The integrated filter modifies the performance-oriented and potentially unsafe control input $\delta^{*}(t)$ to ensure that the actual angle of attack $\alpha(t)$ remains within the safe operational bounds $\alpha_{\text{crit,min}}(t)$ and $\alpha_{\text{crit,max}}(t)$ at all times, preventing violations of stall and structural load constraints. The safety filter optimally adjusts the control input $\delta^{*}(t)$ by solving:
$\min_{\delta,\dot{\delta}}\begin{aligned} &\lVert\delta(t)-\delta^{*}(t)\rVert_{2}^{2}\end{aligned}$ (32)
s.t.
${\psi}_{1,2}(\alpha(t))\geq 0,$ (33)
${\psi}_{2,2}(\alpha(t))\geq 0$ (34)
$\delta_{min}\leq\delta(t)\leq\delta_{max}$ (35)
$\dot{\delta}_{min}\leq\dot{\delta}(t)\leq\dot{\delta}_{max}.$ (36)
Here, $\delta^{*}(t)$ is the desired but potentially unsafe control input, and $\delta(t)$ is the optimized safe input signal passed to the plant. The objective function ensures that $\alpha_{r}(t)$ remains as close as possible to $\alpha_{r}^{*}(t)$, minimizing unnecessary modification while enforcing safety. The HOCBF-derived safety constraints ${\psi}_{1,2}(\alpha(t))\geq 0$ and ${\psi}_{1,2}(\alpha(t))\geq 0$ enforce the upper and lower angle-of-attack limits, ensuring $\alpha$ remains within $\alpha_{\text{crit,min}}$ and $\alpha_{\text{crit,max}}$. The final two inequality constraints make sure that $\delta(t)$ stays within the admissible set of control inputs with respect to magnitude and rate limits.
Remark 3
In the integrated QP-based safety filter formulation, the upper and lower bounds on $\alpha(t)$ are enforced via two distinct HOCBF-derived inequality constraints. In general, this setup could lead to feasibility issues in the QP, especially when both constraints become simultaneously active or conflicting. However, due to the structure of the considered FEP problem, this conflict is typically avoided in practice. Specifically, the system either approaches the upper or the lower bound of the operational $\alpha(t)$ limits depending on the direction of motion—i.e., whether $\alpha(t)$ is increasing or decreasing. As a result, only one of the two HOCBF constraints is usually active at a given time, which significantly reduces the risk of mutual constraint conflict. Nevertheless, when considering input magnitude and rate constraints, additional care must be taken to ensure feasibility. The tuning of the HOCBF parameters becomes crucial to meet the necessary feasibility conditions, particularly when actuator saturation or limited control authority is present. As discussed in [28], enforcing multiple constraints via HOCBFs in the presence of input bounds and disturbances can inherently lead to infeasibility. To address this, feasibility guarantees can be ensured through careful selection of the parameters for the barrier functions. These conditions are further analyzed in the following section, which outlines sufficient tuning strategies to maintain forward invariance while preserving the feasibility of the integrated safety filter.
4.2.1 Derivation of CBF-based Safety Constraints
In the following, we formally derive the safety inequality constraints used in the QP-based safety filter to ensure adherence to the prescribed angle-of-attack limits. While the QP framework has already been introduced in Eq. (32), the explicit formulation of the constraints has not yet been discussed in detail. To address this, we construct HOCBF constraints that enforce the operational bounds on $\alpha$ dynamically, ensuring that the system remains within the permissible flight envelope at all times, as defined in Eq. (31). To do so, we introduce the candidate barrier functions for the upper and lower angle-of-attack constraints:
$h_{1}(\alpha(t))=\alpha_{\text{crit,max}}(t)-\alpha(t),$ (37)
$h_{2}(\alpha(t))=\alpha(t)-\alpha_{\text{crit,min}}(t).$ (38)
These functions define the safe regions:
$S_{1}(t)=\{\alpha(t)\in\mathbb{R}:h_{1}(\alpha(t))\geq 0\},$ (39)
$S_{2}(t)=\{\alpha(t)\in\mathbb{R}:h_{2}(\alpha(t))\geq 0\}.$ (40)
We begin by deriving the constraint for the upper bound on $\alpha$ and subsequently extend the formulation to the lower bound in an analogous manner based on the defined candidate from Eq.(37) for the associated safety set of Eq. (38). As outlined before, it can be assumed $\alpha_{\text{crit,min}}(t)$ and $\alpha_{\text{crit,max}}(t)$ vary slowly, suggesting $\dot{\alpha}_{\text{crit,max}}(t)\approx 0$ and $\dot{\alpha}_{\text{crit,min}}(t)\approx 0$. Since $\alpha(t)$ exhibits second-order dynamics, we construct a second-order HOCBF by recursively defining:
$\psi_{1,0}(\alpha(t))$ $=h_{1}(\alpha(t)),$ (41)
with $\gamma_{1,1}$ and $\gamma_{1,2}$ being positive constants. As later used in Eq. (49), $\psi_{2,2}$ for the HOCBF constraint of the lower state bound on $\alpha$ can be derived very similar to the one of $\psi_{1,2}$. The variables $\dot{\alpha}(t)$ and $\ddot{\alpha}(t)$ are higher order state information of the considered dynamics. However, since no functional relationship on $\ddot{\alpha}(t)$ is directly available, we propose to estimate the needed higher-order information of the vehicle dynamics based on known dynamics. To do so, we first consider again Eq. (21):
With integrating Eq. (21) and Eq. (22) back into the equation, we obtain:
[…]
5 Numerical Results
@@
The missile pitch dynamics model utilized in this study is an adjusted version of the dynamical model shared in [31]. The model is a reduced-order version of a six-degree-of-freedom model, where non-relevant dynamics have been removed. The integrated actuator dynamics are modeled as first-order systems with predefined control surface deflection and rate limits. The relevant environmental, aerodynamic coefficients, and vehicle parameters used in this study are provided in Table 2.
@@
Considered Constraints
Angle of attack limits $\alpha$$\pm 15^{\circ}$
Normal load factor limits $n_{z}$$\pm 10$

### v2

@@ 1 Introduction
Recent advancements in control theory offer an alternative perspective on designing FEP systems, framing the problem as a closed-loop state-constrained control problem. In such a framework, the controllers not only track desired set points but also dynamically enforce safety constraints throughout the system’s operation, both during steady-state and transient phases. Control Barrier Functions (CBFs) have emerged as a powerful tool within this class of safety-critical control methods. CBFs guarantee the forward invariance of a defined safe set, providing formal assurances that the system remains within prescribed limits under all operational conditions [11]. The optimization-based nature of CBFs allows for seamless integration of a wide range of constraints, including physical limitations such as load factor and stall prevention, as well as operational constraints like altitude limits and obstacle avoidance [12, 13]. A significant advantage of CBF-based methods is their ability to augment existing control architectures. Rather than requiring a complete redesign, CBFs can function as an additional safety layer, leveraging the existing closed-loop dynamics as part of the optimization problem [14, 15]. This compatibility facilitates their implementation in modern aerospace systems, while the real-time adaptability of CBFs enables optimal performance under varying conditions, disturbances, and uncertainties [16, 11, 17]. Despite their potential, the application of CBFs to FEP remains relatively underexplored in the flight control community.
@@ 2.1 The Mathematical Concept of Control Barrier Functions
The above definitions lead to a less restrictive version of Eq. (7) as it weakens the requirement to that in Eq. (8). The derived concept is illustrated in Fig. 1. By definition, a valid barrier function $h(x(t))$ is defined as positive within the safe set $S$. To provide safety guarantees and forward invariance within $S$, the condition on the boundary of the safety set needs to be $\dot{h}(x(t))\geq 0$. When this condition is satisfied, it can be proven that the system remains safe.
@@
Since the relevant nonlinear dynamics of the flight systems considered in this work are of a relative degree greater than one, we extend the discussion of CBFs with the notion of High Order Barrier Functions (HOBFs) and High Order Control Barrier Functions (HOCBFs), as proposed in [22]. The control-theoretical construct of CBFs introduced here, along with their extension to HOCBFs, provide a unified methodology for enforcing state constraints derived from operational limits, such as aerodynamic, structural, and performance constraints, which are critical for ensuring safe flight with complex and nonlinear behaviors. By establishing these theoretical foundations, the framework sets the stage for its application to practical systems, as demonstrated in this work through its integration into an FEP scheme. we consider a time-varying function to define an invariant set for system (1). For a $d$-th order differentiable function $h:\mathbb{R}^{n}\times[t_{0},\infty)\to\mathbb{R}$ (where $t_{0}$ denotes the initial time), we define a series of functions $\psi_{0},\psi_{1},\ldots,\psi_{d}$ recursively as:
@@
When a control-affine system of the form (9) is considered, HOBFs can be extended to High Order Control Barrier Functions (HOCBFs) by incorporating control inputs.
@@
This framework introduces a systematic approach to ensure system safety by directly incorporating control inputs into the analysis. Unlike conventional reference filtering methods, HOCBFs enforce safety constraints as an inherent property of the system dynamics, ensuring that operational limits—such as those on angle of attack, load factor, or aerodynamic stability—are satisfied at all times. This makes HOCBFs particularly well-suited for applications in aerospace systems, where higher-order dynamics must be carefully managed to guarantee safe and efficient flight operations.
@@ 2.2 Considered Nonlinear Missile Dynamics
Under the assumption that the roll motion is rapidly stabilized, the missile motions can be decoupled into two perpendicular and identical channels (i.e., pitch and yaw). The missile pitch dynamics model utilized in this study is an adjusted version of the dynamical model shared in [24]. The model is a reduced-order version of a six-degree-of-freedom model, where non-relevant dynamics have been removed.
For the here presented work, the pitch motion is considered as follows:
@@
$\dot{q}(t)$ $=\frac{QSd}{I_{yy}}\left(C_{M,\alpha}(M)\alpha(t)+C_{M,q}(M)\frac{d}{2V}q(t)+C_{M,\delta}(M)\delta(t)\right),$ (22)
$n_{z}(t)$ $=\frac{QS}{mg}\left(C_{Z,\alpha}(M)\alpha(t)+C_{Z,\delta}(M)\delta(t)\right),$ (23)
@@
The integrated actuator dynamics are modeled as first-order systems with predefined control surface deflection and rate limits. The relevant environmental, aerodynamic coefficients, and vehicle parameters used for the flight dynamical modeling in this study are provided in Table 1.
@@
Control Input Constraints
@@
3 The Flight Envelope Protection Problem
For flight systems, aerodynamics are the primary driver of the vehicle’s dynamics, governing its performance and stability during flight. Aerodynamic forces and moments arise from the interaction between the vehicle and the surrounding air, affected by control surfaces and incidence angles such as the angle of attack $\alpha(t)$ and sideslip angle $\beta(t)$. A flight envelope, for any flight vehicle, can be effectively defined and graphically illustrated using a V-n diagram, as shown in Fig. 3. This diagram illustrates the upper and lower load limits corresponding to different airspeeds. These limits provide critical insights into the vehicle’s operational constraints and form the basis for understanding load restrictions.
$x(t_{0})\in S(t)\implies x(t)\in S(t),\quad\forall t\geq t_{0},$ (26)
4 Nonlinear Flight Control System Design with Guaranteed FEP
This section presents the design of a nonlinear flight control system integrated with a FEP safety filter. The overall architecture of the established control system is illustrated in a block diagram in Fig. 4.
Figure 4: Block diagram of the integrated control architecture with a FEP safety filter.
First, a cascaded nonlinear attitude flight controller is designed in Sect. 4.1 to regulate the nonlinear longitudinal missile dynamics. Following this, a CBF-based safety filter is proposed in Sect. 4.2 to ensure adherence to predefined safety constraints, effectively maintaining the missile system’s operation, defined in Sect 2.2, within the defined flight envelope while minimally influencing control performance when the dynamics are far from the envelope boundaries.
4.1 Nonlinear Attitude Controller Design
$\dot{\alpha}_{c}(t)=K_{\alpha}(\alpha_{r}(t)-\alpha(t)),$ (27)
$\dot{q}_{c}(t)=K_{q}(q_{c}(t)-q(t)),$ (29)
Controller Gain Conservative Aggressive
$K_{\alpha}$ 4.00 36.53
$K_{q}$ 1.2 12.22
Remark 2
It should be noted that, as will be shown in the following section, the potentially unsafe control surface deflection $\delta^{*}(t)$ will be adjusted to ensure that the flight vehicle remains within its designated flight envelope. However, this adjustment comes at the cost of deviating from the reference command $\alpha_{r}(t)$. This deviation can introduce challenges, particularly when integral controllers are employed within the integrated flight control system to ensure adequate tracking performance. In such cases, integrator windup may occur, leading to undesirable control behavior and degraded system performance. To mitigate these effects, an windup protection approach must be implemented to prevent excessive error accumulation and maintain stable control performance.
4.2 CBF-Based Nonlinear Safety Filter Design
Fig. 7 illustrates the qualitative relationship between airspeed $V$ and the permissible limits on the angle of attack $\alpha(t)$, which must not be violated to ensure safe operation. The dashed-dotted black lines represent structural constraints, the dotted black lines indicate stall boundaries, and the solid red lines delineate the combined overall operational limits of $\alpha(t)$ as a function of airspeed. The operational limits on $\alpha(t)$ imposed by structural load constraints can be derived from Eq. (23), considering the strictly defined structural load factor limits $n_{z,\text{max,structural}}$ and $n_{z,\text{min,structural}}$ the equation can be inverted to obtain the corresponding maximum and minimum permissible angles of attack $\alpha_{\text{max,structural}}$ and $\alpha_{\text{min,structural}}$. The structural constraints thereby impose an inverse-square dependency of $\alpha(t)$ on $V$, owing to the quadratic dependence of $Q$ on $V$. In contrast, stall limits are typically assumed to be independent of airspeed, meaning that the maximum and minimum allowable angles of attack due to stall, $\alpha_{\text{max,stall}}$ and $\alpha_{\text{min,stall}}$, remain constant across the flight envelope. As a result, the operational bounds on $\alpha(t)$ are shaped by the interplay between structural and stall constraints. At lower airspeeds, stall limits dominate the constraints on $\alpha(t)$. In contrast, at higher airspeeds, the range of feasible $\alpha(t)$ values is increasingly restricted by structural considerations, as illustrated in Fig. 7 with solid red lines. By considering the described two origins of constraints on $\alpha(t)$, the overall operational limits can be formally expressed as:
where:
$\alpha_{\text{crit,min}}(t)$ $=\max(\alpha_{\text{min,stall}}(t),\alpha_{\text{min,structural}}(t)).$
To dynamically enforce the derived angle-of-attack constraints, a HOCBF-based safety filter is implemented. The integrated filter modifies the performance-oriented and potentially unsafe control input $\delta^{*}(t)$ to ensure that the actual angle of attack $\alpha(t)$ remains within the safe operational bounds $\alpha_{\text{crit,min}}(t)$ and $\alpha_{\text{crit,max}}(t)$ at all times, preventing violations of stall and structural load constraints. The safety filter optimally adjusts the control input $\delta^{*}(t)$ by solving:
$\min_{\delta,\dot{\delta}}\begin{aligned} &\lVert\delta(t)-\delta^{*}(t)\rVert_{2}^{2}\end{aligned}$
s.t.
${\psi}_{1,2}(\alpha(t))\geq 0,$
${\psi}_{2,2}(\alpha(t))\geq 0,$ (32)
$\delta_{min}\leq\delta(t)\leq\delta_{max},$
$\dot{\delta}_{min}\leq\dot{\delta}(t)\leq\dot{\delta}_{max}.$
Here, $\delta^{*}(t)$ is the desired but potentially unsafe control input, and $\delta(t)$ is the optimized safe input signal passed to the plant. The objective function ensures that $\alpha_{r}(t)$ remains as close as possible to $\alpha_{r}^{*}(t)$, minimizing unnecessary modification while enforcing safety. The HOCBF-derived safety constraints ${\psi}_{1,2}(\alpha(t))\geq 0$ and ${\psi}_{1,2}(\alpha(t))\geq 0$ enforce the upper and lower angle-of-attack limits, ensuring $\alpha(t)$ remains within $\alpha_{\text{crit,min}}$ and $\alpha_{\text{crit,max}}$. The final two inequality constraints make sure that $\delta(t)$ stays within the admissible set of control inputs with respect to magnitude and rate limits.
Remark 3
In the integrated QP-based safety filter formulation, the upper and lower bounds on $\alpha(t)$ are enforced via two distinct HOCBF-derived inequality constraints. In general, this setup could lead to feasibility issues in the QP, especially when both constraints become simultaneously active or conflicting. However, due to the structure of the considered FEP problem, this conflict is typically avoided in practice. Specifically, the system either approaches the upper or the lower bound of the operational $\alpha(t)$ limits depending on the direction of motion—i.e., whether $\alpha(t)$ is increasing or decreasing. As a result, only one of the two HOCBF constraints is usually active at a given time, which significantly reduces the risk of mutual constraint conflict. Nevertheless, additional care must be taken to ensure feasibility when input constraints are present. As discussed in [29], feasibility guarantees can be ensured through careful selection of the design parameters for the integrated CBFs. This is further discussed in the following part of this section, which outlines a tuning strategy to maintain forward invariance while preserving the feasibility of the integrated safety filter.
4.2.1 Derivation of CBF-based Safety Constraints
While the QP framework has already been introduced in Eq. (32), the explicit formulation of the constraints has not yet been discussed in detail. To address this, we construct HOCBF constraints that enforce the operational bounds on $\alpha(t)$ dynamically, ensuring that the system remains within the permissible flight envelope at all times, as defined in Eq. (31). To do so, we introduce the candidate barrier functions for the upper and lower angle-of-attack constraints:
$h_{1}(\alpha(t))=\alpha_{\text{crit,max}}(t)-\alpha(t),$ (33)
$h_{2}(\alpha(t))=\alpha(t)-\alpha_{\text{crit,min}}(t).$ (34)
These functions define the safe regions:
$S_{1}(t)=\{\alpha(t)\in\mathbb{R}:h_{1}(\alpha(t))\geq 0\},$ (35)
$S_{2}(t)=\{\alpha(t)\in\mathbb{R}:h_{2}(\alpha(t))\geq 0\}.$ (36)
We begin by deriving the constraint for the upper bound on $\alpha(t)$ and subsequently extend the formulation to the lower bound in an analogous manner based on the defined candidate from Eq.(33) for the associated safety set of Eq. (34). As outlined before, it can be assumed $\alpha_{\text{crit,min}}(t)$ and $\alpha_{\text{crit,max}}(t)$ vary slowly, suggesting $\dot{\alpha}_{\text{crit,max}}(t)\approx 0$ and $\dot{\alpha}_{\text{crit,min}}(t)\approx 0$. Since $\alpha(t)$ exhibits second-order dynamics, we construct a second-order HOCBF by recursively defining:
$\psi_{1,0}(\alpha(t))$ $=h_{1}(\alpha(t)),$ (37)
with $\gamma_{1,1}$ and $\gamma_{1,2}$ being positive constants. As later used in Eq. (45), $\psi_{2,2}$ for the HOCBF constraint of the lower state bound on $\alpha(t)$ can be derived very similar to the one of $\psi_{1,2}$. The variables $\dot{\alpha}(t)$ and $\ddot{\alpha}(t)$ are higher order state information of the considered dynamics. However, since no functional relationship on $\ddot{\alpha}(t)$ is directly available, we propose to estimate the needed higher-order information of the vehicle dynamics based on known dynamics. To do so, we first consider again Eq. (21):
With integrating Eq. (42) back into the Eq. (41), we obtain:
[…]
5 Numerical Results
A simulation-based analysis was carried out to examine the proposed FEP methodology. The flight dynamics model presented in Sect. 2.2 was implemented in a MATLAB/Simulink environment.